OpenStudy (anonymous):

Will Fan and Medal What is the solution to the equation? 2(x – 7) = 10x + 18 A. x=4 B. x=15/8 C. x=-4 D. X=-15/8

OpenStudy (muzzack):

okay, there is a lil trick i know

OpenStudy (anonymous):

I say B

OpenStudy (muzzack):

replace the x with each of them

OpenStudy (muzzack):

try B first because @jbwing said it

OpenStudy (anonymous):

YASSSSS lol @Muzzack

OpenStudy (muzzack):

lol :)

OpenStudy (muzzack):

x_O

OpenStudy (anonymous):

are u guys sure its be this is the semester assessment it cant be rong

OpenStudy (muzzack):

no im not saying it B im saying try them out and if they give you for ex. 5=5 it is right :)

OpenStudy (yanasidlinskiy):

2(x – 7) = 10x + 18 2x-14 = 10x + 18 ------------variables on both sides -2x -2x -14 = 8x + 18 -18 -18 -32 = 8x -------- divide by 8 to get rid of x -4 = x

OpenStudy (yanasidlinskiy):

Any questions?

OpenStudy (anonymous):

@YanaSidlinskiy thank you

OpenStudy (anonymous):

yes im about to post it @YanaSidlinskiy

OpenStudy (solomonzelman):

2(x – 7) = 10x + 18 2(x – 7) = 2(x + 9) x-7=x+9 add 7 to both sides x=x+18 this is no solution, @ayeeitsme do you understand why ?

OpenStudy (yanasidlinskiy):

Your Welcome!:)